OpenStudy (anonymous):

MEDAL!! Solve equation, giving a general formula for all the solutions. 2sin^2 x-5sinx-3=0

OpenStudy (anonymous):

I know the answers are 7pi/6 and 11pi/6, but I don't know the general formula

OpenStudy (anonymous):

@ksaimouli

OpenStudy (anonymous):

is it x=pi/6 +kpi

OpenStudy (anonymous):

@ikram002p is this correct?

OpenStudy (ikram002p):

limme check

OpenStudy (ikram002p):

2sin^2 x-5sinx-3=0 (2sin x +1) (sin x-3)=0 2 sin x +1 = 0 sin x =1/2 x= sin^-1 1/2=pi/6 + kpi ur correct

OpenStudy (anonymous):

so that is the general formula!!??

OpenStudy (ikram002p):

pi/6 + k2pi